### v1

# Fermi-liquid transport beyond the upper critical field in superconducting La2PrNi2O7 thin films

Fermi-liquid transport beyond the upper critical field in superconducting La2PrNi2O7 thin films
Yu-Te Hsu***These authors contributed equally to this work.ythsu@phys.nthu.edu.tw Department of Physics, National Tsing Hua University, Hsinchu 300044, Taiwan Department of Materials Science and Engineering, National Tsing Hua University, Hsinchu 300044, Taiwan Yidi Liu Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Physics, Stanford University, Stanford, CA 94305, United States Yoshimitsu Kohama Institute for Solid State Physics, The University of Tokyo, Kashiwa, Chiba 277-8581, Japan Tommy Kotte Dresden High Magnetic Field Laboratory (HLD-EMFL), Helmholtz-Zentrum Dresden-Rossendorf, Dresden 01328, Germany Vikash Sharma Institute for Solid State Physics, The University of Tokyo, Kashiwa, Chiba 277-8581, Japan Yaoju Tarn Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States Yijun Yuyuyijun@stanford.edu Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States Harold Y. Hwang Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States
Unconventional superconductivity typically emerges out of a strongly correlated normal state, manifesting as a Fermi liquid with highly enhanced effective mass or a strange metal with $T$-linear resistivity in the zero-temperature limit. In Ruddlesden-Popper bilayer nickelates $R_{3}$i2O7, superconductivity with a critical temperature ($T_{\rm c}$) exceeding 80 and 40 K has been respectively realised in bulk crystals under high pressure and thin films under compressive strain. These advancements create new materials platforms to study the nature of high-$T_{\rm c}$ superconductivity, calling for the characterisation of fundamental normal-state and superconducting parameters therein. Here we report detailed magnetotransport experiments on superconducting La2PrNi2O7 (LPNO) thin films under pulsed magnetic fields up to 64 T and access the normal-state behaviour over a wide temperature range between 1.5 and 300 K. We find that the normal state of LPNO exhibits the hallmarks of Fermi liquid transport, including $T^{2}$ temperature dependence of resistivity and Hall angle, and $H^{2}$ magnetoresistance obeying Kohler scaling. Using the empirical Kadowaki-Woods ratio relating the transport coefficient and electronic specific heat, we estimate a quasiparticle effective mass $m^{*}/m_{e}\simeq 10$ in PLNO, thereby revealing the highly renormalized Fermi liquid state which hosts the high-temperature nickelate superconductivity. Our results demonstrate that LPNO follows the same $T_{\rm c}/T_{\rm F}^{*}$ scaling observed across a wide variety of strongly correlated superconductors and provide crucial constraints for a viable model for superconductivity in bilayer nickelates.
Introduction
The recent discovery of superconductivity at 80 K in (La,Pr)3Ni2O7 under high pressure makes the bilayer Ruddlesden-Popper (RP) nickelates the latest material system exhibiting high-temperature superconductivity (HTSC) [1, 2]. Despite intense interests in the nature of superconductivity in RP bilayer nickelates [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], key parameters characterising the normal and superconducting state are still missing. This is in large part due to the high-pressure condition required to induce superconductivity, which limits the type of experiments that can be performed to primarily transport under static magnetic field and x-ray diffraction. The subsequent realisation of superconductivity in epitaxially strained (La,Pr)3Ni2O7 thin films with a critical temperature $T_{\rm c}\approx$ 40 K (refs. [13, 14, 15]) creates a new material platform to which a broader selection of experimental probes can be employed at ambient pressure, including photoemission [16, 17], x-ray spectroscopy [18], and electron energy-loss spectroscopy [19]. Moreover, the availability of superconducting (La,Pr)3Ni2O7 thin films enables a straightforward implementation of pulsed high-field magnetotransport experiment, which can reveal a wealth of information on its superconducting and normal state, including upper critical field ($H_{\rm c2}$), carrier concentration ($n_{\rm H}$), and functional form of low-temperature normal-state resistivity $\rho_{\rm n}(T)$.
In superconducting (La,Pr)3Ni2O7 above $T_{\rm c}$, $\rho_{\rm xx}(T)$ in bulk crystal and thin film shows an intriguing dichotomy: a putative strange-metal $T$-linear resistivity is found in bulk crystals [1, 20, 2] whereas an apparent Fermi-liquid $T$-quadratic resistivity is found in thin films[14, 15]. As the identification of the non-superconducting electronic ground state is critical to understanding the mechanism of electron pairing and the formation of superconducting condensate, it is essential to establish the form of $\rho_{\rm n}(T)$ below $T_{\rm c}$ in the bilayer nickelates. It has been recently found that a partial substitution of La by Pr in (La,Pr)3Ni2O7 favours the formation of bilayer structure [2, 14, 15], leading to a higher crystallinity as reflected by a residual resistivity ratio as high as 8 (ref. [15]), therefore we choose La2PrNi2O7 (LPNO) thin film as the model platform to study the magnetotransport behaviour of bilayer superconducting nickelates.
Prior magnetotransport experiments on LPNO thin film estimate that $\mu_{0}H_{\rm c2}$ in the $T=0$ limit is well above 50 T [13, 14, 15], necessitating the use of pulsed field conditions to access the low-$T$ normal state. Here, we report detailed magnetotransport experiment on superconducting LPNO thin films under pulsed magnetic fields up to 64 T. We reveal salient characteristics of the normal-state transport in LPNO, including 1) a $T^{2}$ resistivity in the field-induced low-$T$ normal state, 2) a saturating Hall coefficient as $T\rightarrow 0$ and a $T^{2}$ temperature dependence of the Hall angle, and 3) an observation of Kohler scaling in normal-state magnetoresistance. These findings provide strong evidence that the underlying electronic ground state in strained LPNO thin film is a Fermi liquid. Our measurements also reveal that $H_{\rm c2}(T=0)$ in LPNO exhibits an anisotropy with $\gamma=H_{c2,\parallel}/H_{\rm c2,\perp}\approx 2.5$. By estimating the magnitude of key normal-state and superconducting parameters using established approaches, we find that the ratio between $T_{\rm c}$ and effective Fermi temperature $T_{\rm F}^{*}$ in bilayer RP nickelates exhibits a scaling of $T_{\rm c}/T_{\rm F}^{*}\sim 0.01$, as found in a wide variety of strongly-correlated superconductors [21, 22, 23, 24], thereby hinting at a universal underlying principle governing the $T_{\rm c}$ magnitude in unconventional superconductors.
Results
Form of normal-state resistivity below $T_{\rm c}$
Figure 1a shows the in-plane resistivity $\rho_{\rm xx}(H,T)$ of a LPNO thin film in pulsed magnetic field up to 53 T. The field is applied parallel to the crystalline $c$-axis i.e. perpendicular to the film surface. A clear transition from a superconducting to a resistive state can be seen below $T_{\rm c}=41$ K (defined using the 90% $\rho_{\rm n}$ criterion with $\rho_{\rm n}(T)$ obtained by the parallel-resistor fit to resistivity above 50 K). The temperature dependence of resistivity measured in the field-induced normal state at 53 T and extrapolated to 0 T are plotted in Fig. 1b. A prominent feature of $\rho_{\rm xx}(T)$ is the positive curvature as $T\rightarrow 0$ and negative curvature as $T\rightarrow 300$ K. As demonstrated in previous work [15], $\rho_{\rm xx}(T)$ above $T_{\rm c}$ can be well-described using a parallel resistor formula (PRF):
$1/\rho(T)=1/({\rho_{0}+A_{2}T^{2}})+1/\rho_{\rm max},$ (1)
where $\rho_{0}$, $\rho_{\rm max}$, and $A_{2}$ are fitting constants. By suppressing superconductivity with large magnetic field, we find that the field-induced normal-state resistivity below $T_{\rm c}$ tracks the PRF fit very closely and shows a clear saturation towards an apparent $T^{2}$-behaviour as $T\rightarrow 0$ (Fig. 1b inset). The same behaviour is reproduced in a second sample (Supplementary Materials Fig. S1). Furthermore, the extrapolated zero-field resistivity exhibits essentially the same $\rho(T)$ form as $\rho_{\rm xx}(\textrm{53\leavevmode\nobreak\ T})$ (see details for zero-field resistivity extrapolation in Methods), demonstrating that the normal-state $\rho_{\rm xx}(T)$ of LPNO exhibits a $T^{2}$ functional form in the $T=0$ limit, characteristic of a Fermi-liquid ground state.
Figure 1: In-plane resistivity of a superconducting La2PrNi2O7 thin film. a, Magnetoresistivity isotherms $\rho_{\rm xx}(H)$ up to 53.4 T measured at the following temperatures: 100, 80, 60, 50, 45, 40, 35, 30, 25, 20, 15, 10, 7.0, 4.1, and 1.5 K. b, Temperature-dependent resistivity $\rho_{\rm xx}(T)$ at 0 and 53 T. Grey dash line is a fit to the measured zero-field resistivity above between 50 and 300 K using a parallel resistor model: $1/\rho(T)=1/(\rho_{0}+A_{2}T^{2})+1/\rho_{\rm max}$, which finds $\rho_{0}=102\leavevmode\nobreak\ {\rm\mu\Omega\leavevmode\nobreak\ cm}$, $\rho_{\rm max}=778\leavevmode\nobreak\ {\rm\mu\Omega\leavevmode\nobreak\ cm}$, and $A_{2}=8.1\leavevmode\nobreak\ \rm{n\Omega\leavevmode\nobreak\ \leavevmode\nobreak\ cm\leavevmode\nobreak\ K^{-2}}$, respectively. Inset: $\rho_{\rm xx}$ versus $T^{2}$ below 60 K, showing a $\Delta\rho(T)\leavevmode\nobreak\ \propto\leavevmode\nobreak\ T^{2}$ behaviour in the measured 53-T resistivity and extrapolated zero-field resistivity.
Hall effect and magnetoresistance
Figure 2a shows the Hall resistivity $\rho_{\rm yx}(H)$ measured between 10 and 60 K on the same sample as shown in Fig. 1. $H$-linear behaviour in $\rho_{\rm yx}(H)$ (within the noise level) is found in the (field-induced) normal state, consistent with the low-field measurements above $T_{\rm c}$ in La3Ni2O7 and LPNO thin films [13, 15]. The temperature dependence of Hall coefficient $R_{\rm H}(T)$ between 10 and 240 K, plotted in Fig. 2a inset, shows a monotonic increase as $T$ decreases and approaches a constant value of $R_{\rm H}\approx$ 0.3 mm3/C as $T\rightarrow 0$. We estimate the total carrier density using $n_{\rm H}(0)\simeq 1/(eR_{\rm H}(0))=2.1\times 10^{22}\leavevmode\nobreak\ \rm{cm}^{-3}$ and from which a two-dimensional (2D) carrier density $n_{\rm 2D}(0)=n_{\rm H}(0)d=2.1\times 10^{15}\rm\leavevmode\nobreak\ {cm}^{-2}$ per NiO2 bilayer unit ($d=10.3$ is the distance between neighboring NiO2 bilayers [2]). Note that this $n_{\rm H}$ value should be interpreted as a coarse estimate for the inference of $k_{\rm F}$ and possibly not constitute an accurate quantification of the total carrier density, as details of the electronic structure in bilayer nickelates remain to be established.[25, 16] Nevertheless, assuming a free-electron model in two dimensions, we find $k_{\rm F}=1.15\times 10^{10}\leavevmode\nobreak\ \rm{m}^{-1}$, comparable to that of the $\beta$ sheet found by recent photoemission experiments[25, 16].
Figure 2: Hall resistivity, Kohler scaling, and Hall angle in La2PrNi2O7 thin film. a, Hall resistivity isotherms $\rho_{\rm yx}(H)$ measured at indicated temperatures. Dash lines are fits made to the normal-state $\rho_{\rm yx}$ at high fields using $R_{\rm H}=\rho_{\rm yx}/(\mu_{0}H)$. Inset: Hall coefficient versus temperature $R_{\rm H}(T)$. A saturation of $R_{\rm H}\simeq 0.3$ mm3/C is found as $T\rightarrow 0$. Vertical dashed line marks $T_{\rm c}$. b, Normalised magnetoresistance versus magnetic field scaled by zero-field resistivity i.e. $\rho_{\rm xx}/\rho_{\rm xx}(0)$ versus $\mu_{0}H/\rho_{\rm xx}(0)$, known as the Kohler plot. Normal-state data measured at indicated temperatures collapses into a single curve following $\rho_{\rm xx}/\rho_{\rm xx}(0)\leavevmode\nobreak\ \propto\leavevmode\nobreak\ (\mu_{0}H/\rho_{\rm xx}(0))^{2}$ shown by the dashed line. c, Inverse Hall angle $\cot\theta_{\rm H}=\rho_{\rm xx}/\rho_{\rm xx}$ versus $T^{2}$. Filled points correspond to measurements at 53 T and open points correspond to extrapolated 53 T values using Hall resistivity measured at 8 T (i.e. $\rho_{\rm yx}(\rm 53\leavevmode\nobreak\ T)$ = $R_{\rm H}(8\leavevmode\nobreak\ \rm{T})\times(53/8)$). Dashed line is a fit using $\cot\theta_{\rm H}=C+BT^{2}.$
The form of normal-state magnetoresistance (MR = $\Delta\rho_{\rm xx}(H)/\rho_{\rm xx}(0)$) is examined in Fig. 2b. The MR magnitude is small ($\lesssim 4\%$ up to 53 T) and follows a $H^{2}$-behaviour i.e. $\Delta\rho(H)/\rho(0)\leavevmode\nobreak\ \propto\leavevmode\nobreak\ H^{2}$. We find that $\rho_{\rm xx}(H)/\rho_{\rm xx}(0)$ collapse into a single curve when plotted against $\mu_{0}H/\rho_{\rm xx}(0)$, known as the Kohler scaling, which implies that the MR magnitude is dictated by the carrier mean free path $\ell\sim\omega_{\rm c}\tau$ (i.e. the product of cyclotron frequency $\omega_{\rm c}$ and carrier lifetime $\tau$), as expected for a Fermi-liquid state characterised by a single quasiparticle lifetime [26]. The fact that the normal-state MR $\propto\leavevmode\nobreak\ H^{2}$ and obeys Kohler scaling allows us to extract the zero-field normal-state resistivity below $T_{\rm c}$ as shown in Fig. 1b (see Methods for details). The Fermi-liquid nature of magnetotransport in LPNO is further supported by the observation of a negligible MR when the magnetic field is applied parallel to the sample surface i.e. $\mathbf{H}\parallel ab$ (Supplementary Materials Fig. S2) and a $T^{2}$ dependence of the Hall angle: $\cot\theta_{\rm H}=\rho_{\rm xx}/\rho_{\rm yx}\leavevmode\nobreak\ \propto\leavevmode\nobreak\ (\omega_{\rm c}\tau)^{-1}$ (Fig. 2c). These features are in stark contrast to the high-$T_{\rm c}$ cuprates within the strange-metal regime, in which a sizeable MR is found with $\mathbf{H}\parallel ab$ (ref. [27]) and the longitudinal resistivity and Hall angle exhibit distinct $T$-dependence [28], suggesting that the longitudinal and Hall transport are governed by two distinct carrier lifetimes. Overall, we find that the normal-state transport characteristics of superconducting LPNO, including $\rho_{\rm xx}(T)$, MR, Hall effect, and Hall angle, can all be well-described using the standard transport theory for a conventional Fermi-liquid, the key finding of this work.
Upper critical field anisotropy
Next we extract the upper critical field $H_{\rm c2}$ in both $\mathbf{H}\parallel c$ and $\mathbf{H}\parallel ab$ configuration (i.e. $H_{\rm c2,\perp}$ and $H_{\rm c2,\parallel}$, respectively). Due to the considerable transition width in magnetic field, the choice of different extraction criteria leads to different forms of $H_{\rm c2}(T)$ (see Supplementary Materials Fig. S3). We adopt the 90% $\rho_{\rm n}$ criterion, in accordance with previous reports [13, 15], and extract $H_{\rm c2}$ using the condition $\rho_{\rm xx}(\mu_{0}H_{\rm c2})/\rho_{\rm xx}(\rm{53\leavevmode\nobreak\ T)}=0.9$. The $\mu_{0}H_{\rm c2}(T)$ data is fitted using the linearised Ginzburg-Landau form [29]:
$H_{\rm c2,\perp}(T)=\frac{\phi_{0}}{2\pi\xi_{ab}^{2}}\left(1-\frac{T}{T_{\rm c}}\right),$ (2)
$H_{\rm c2,\parallel}(T)=\frac{\sqrt{12}\phi_{0}}{2\pi\xi_{ab}(0)d_{\rm sc}}\left(1-\frac{T}{T_{\rm c}}\right)^{\frac{1}{2}},$ (3)
where $\phi_{0}$ is the flux quantum, $\xi_{ab}(0)$ is the in-plane coherence length at zero temperature, and $d_{\rm sc}$ is the effective superconducting thickness. A highly anisotropic $H_{\rm c2}$ is found with $\mu_{0}H_{\rm c2,\perp}(0)=(42.8\pm 0.4)$ T and $\mu_{0}H_{\rm c2,\parallel}(0)=(106\pm 1)$ T, and a corresponding anisotropic factor $\gamma=H_{\rm c2,\parallel}/H_{\rm c2,\perp}=2.48$. Furthermore, we find the in-plane coherence length $\xi_{ab}(0)=(2.76\pm 0.01)$ nm and a superconducting thickness $d_{\rm sc}=(3.88\pm 0.02)$ nm, comparable to the film thickness of 5 nm and suggesting a bulk nature of superconductivity.
Comparing to the previously reported values for La3Ni2O7 and LPNO thin films [13, 15], our extracted $\mu_{0}H_{\rm c2,\perp}(0)$ is significantly lower. We note that in previous reports, a constant resistivity value at a temperature slightly above $T_{\rm c}$ (e.g. 50 K) is used to define the 90% $\rho_{\rm n}$ criterion for $\rho_{\rm xx}(H,T)$ measured at $H\ll H_{\rm c2}$. In LPNO, since the normal-state $\rho_{\rm xx}(T)$ follows a $T^{2}$ behaviour and the $T_{\rm c}\approx$ 40 K is relatively high, the 50-K resistivity value considerably overestimates the true normal-state resistivity at $T\lesssim T_{\rm c}$, thus leading to an overestimation of $H_{\rm c2}(T\lesssim T_{\rm c})$ and $H_{\rm c2}(0)$. It is also likely that there exists a sample variations in $H_{\rm c2}$ (see Supplementary Materials Fig. S1), whose underlying cause requires further investigation. Our current experiment on a LPNO film with an $H_{\rm c2,\perp}$ directly accessible within 53 T provides an accurate quantification of its $\mu_{0}H_{\rm c2,\perp}(0)$, and allows a crude estimate of the superconducting gap magnitude $\Delta_{0}\simeq 9.9$ meV via $\xi_{ab}(0)=\hbar v_{\rm F}/(\pi\Delta_{0})$. We caution here that this $\Delta_{0}$ value should be interpreted as an order-of-magnitude estimate as $v_{\rm F}$ used here is obtained using estimates of $k_{\rm F}$ and $m^{*}$, both of which are associated a sizeable uncertainty, and the level of disorder is assumed to play only a minor role in the extraction of $T_{\rm c}$ and $H_{\rm c2}$ via resistivity measurements.
Figure 3: Upper critical fields of La2PrNi2O7 thin film. $\mu_{0}H_{\rm c2}$ are extracted using the criterion of $\rho_{\rm xx}(\mu_{0}H_{\rm c2})/\rho_{\rm xx}$(53 T) = 0.9. Filled points correspond to the configuration that the magnetic field is applied along the film surface normal (i.e. $\mathbf{H}\parallel c$) and open points the configuration that field applied parallel to the film surface (i.e. $\mathbf{H}\parallel ab$). Dash and dotted lines are fits made using the linearised Ginzburg-Landau formulae (see main text). Inset shows $\mu_{0}H_{c2}$ measured at 35 K with the field oriented close to the in-plane configuration (i.e. $\theta\approx 90^{\circ}$). Solid line is a fit made using the two-dimensional Tinkham model[30]: $\left(\frac{H_{\rm c2}\sin\theta}{H_{\rm c2,\parallel}}\right)^{2}+\left|\frac{H_{\rm c2}\cos\theta}{H_{\rm c2,\perp}}\right|=1$.
Estimate of effective mass
It is widely known that the prefactor of $T^{2}$-resistivity and electronic specific heat $\gamma_{0}$ for a variety of correlated electron systems exhibits an empirical relationship[31, 32]:
$A_{2}/\gamma_{0}^{2}\sim 10^{5}\leavevmode\nobreak\ \mu\Omega\leavevmode\nobreak\ \rm{cm\leavevmode\nobreak\ mJ^{-2}\leavevmode\nobreak\ mol^{-2}\leavevmode\nobreak\ K^{-6}}$, known as the Kadowaki-Woods ratio. The high-pressure or epitaxial strain condition required for superconductivity in the bilayer nickelates currently precludes the extraction of $\gamma_{0}$ via direct calorimetry measurements. Alternatively, by assuming the empirical Kadowaki-Woods ratio of $10^{5}\leavevmode\nobreak\ \mu\Omega\leavevmode\nobreak\ \textrm{cm}\leavevmode\nobreak\ \textrm{mJ}^{-2}\leavevmode\nobreak\ \textrm{mol}^{-2}\leavevmode\nobreak\ \textrm{K}^{-6}$ and using the measured $A_{2}=(8.1\pm 1.2)\leavevmode\nobreak\ \textrm{n}\Omega\leavevmode\nobreak\ \textrm{cm}\leavevmode\nobreak\ \textrm{K}^{-2}$, we infer a $\gamma_{0}=(28\pm 4)\leavevmode\nobreak\ \textrm{mJ}\leavevmode\nobreak\ \textrm{mol}^{-1}\leavevmode\nobreak\ \textrm{K}^{-2}$ for the LPNO film studied in this work. We further estimate the quasiparticle effective mass using[33]:
$\gamma_{0}=\gamma\prime\sum_{i}m_{i}^{*}/m_{e},$ (4)
[…]
Table 1: Normal-state and superconducting parameters for La2PrNi2O7 thin film discussed in the main text. Note that the $\Delta_{0}$ value presented here represents an order-of-magnitude estimate (see main text for discussion).
[…]
$8.1\pm 1.2$ $2.1\pm 0.2$ $2.76\pm 0.01$ $10.3\pm 1.5$ 9.9
Discussion
It has been noted that, for a wide variety of unconventional superconductors, the ratio between $T_{\rm c}$ and $T_{\rm F}^{*}$ share a very similar magnitude i.e. $T_{\rm c}/T_{\rm F}^{*}\sim 0.01$ refs. [21, 24, 22, 23]). A survey of $T_{\rm c}$ versus $T_{\rm F}^{*}$ for notable superconductors, including high-$T_{\rm c}$ cuprates[21], iron-based superconductors[22], heavy fermion materials[23], and carbon-based superconductors[24], is shown in Fig. 4. We find that for LPNO the empirical relationship $T_{\rm c}/T_{\rm F}^{*}\sim 0.01$ also holds, suggesting a strongly correlated nature of superconductivity in the bilayer RP nickelates and the magnitude of $T_{\rm c}$ is governed by a universal underlying principle independent of material-specific details.
Figure 4: Superconducting critical temperature $T_{\rm c}$ versus effective Fermi temperature $T_{\rm F}^{*}$ for strongly correlated superconductors. The dotted, dotted-dash, and dashed lines indicate $T_{\rm c}/T_{\rm F}^{*}$ = 0.01, 0.05, and 0.1, respectively. $T_{\rm F}^{*}$ values are extracted assuming a quadratic energy dispersion $E_{\rm F}=\hbar^{2}k_{\rm F}^{2}/(2m^{*})$ and using experimental data of effective mass and carrier density inferred from specific heat, Hall effect, and quantum oscillation measurements (refs. [21, 24, 22, 23] and references therein) or from penetration depth measurements using $E_{\rm F}=(\hbar^{2}/2)(3\pi^{2})^{2/3}n_{\rm s}^{2/3}/m^{*}$ (ref. [38]). The series of points in gradient blue illustrates the effect of hole doping on $T_{\rm c}$ and $T_{\rm F}^{*}$ of La2-xSrxCuO4 (LSCO) with the darker colour indicating a higher doping level up to $x=0.21$ ($T_{\rm c}=27$ K). Error bars for the cuprate materials indicate the variations in $T_{\rm c}$ and $T_{\rm F}^{*}$ with carrier dopings, and for LPNO the variations in $T^{*}_{\rm F}$ inferred from 0-T resistivity measured on multiple samples[15]. MATBG: magic-angle twisted bilayer graphene; TMTSF: tetramethyltetraselenafulvalene; BEDT: bisethylenedithiol; Ba122: BaFe2(As1-xPx)2; STO: SrTiO3; YBCO: YBa2Cu3O6+x; Tl2201: Tl2Ba2CuO6+δ; Tl2223: Tl2Ba2Ca2Cu3O10+δ; Bi2223: Bi2Sr2Ca2Cu3O10+δ; Hg2223: Hg2Ba2Ca2Cu3O10+δ;
[…]
Another possibility to account for the contrasting $\rho(T)$ is the qualitative different effect between hydrostatic pressure and epitaxial strain, leading to an opposite change in $c$-axis length and possibly critical distinction in the underlying electronic structure. It has been demonstrated that in epitaxial thin film of La2-xSrxCuO4 grown on LaSrAlO4 near optimal doping ($p=0.15$), $\rho(T)$ evolves from being $T$-linear above $T_{\rm c}$ to being $T$-superlinear as the film thickness reduces from 90 to 3 nm [43], accompanied by a reduction in residual resistivity ratio (RRR=$\rho(\textrm{300\leavevmode\nobreak\ K})/\rho_{0}$) from 6 to 2. We note that in the case of La3Ni2O7 and LPNO, the bulk crystals showing a $T$-linear resistivity have RRR values of 2.0–2.5 (refs.[1, 20, 2]), whereas our thin films show RRR $\approx 4$ (Fig. 1 and Supplementary Materials, Fig. S1). It is thus unlikely that the deviation from a $T$-linear resistivity in LPNO is caused by an enhanced level of disorder, and that the distinction in $\rho(T)$ functional form between that found in bulk crystals and thin films reflects a difference in hole doping and/or strain anisotropy.
Lastly, we comment on the surprising fact that HTSC emerges out of a Fermi-liquid ground state in LPNO. While a $T^{2}$-resistivity is expected for a conventional metallic state as $T\rightarrow 0$, it is generally not expected that such resistivity functional form will hold at high temperatures; instead, $\rho(T)$ in conventional metals typically exhibits a $T$-linear behaviour at intermediate temperature range due to dominant electron-phonon coupling. In the current case of thin-film LPNO, in no temperature range up to 300 K is a $T$-linear resistivity found, implying that the electron-electron interaction in the bilayer nickelates is so dominant that the electron-phonon coupling becomes essentially negligible. This dominant electron-electron coupling is likely to be relevant for understanding the formation of density wave orders in bilayer nickelates [44, 45, 46]. If LPNO is to be mapped to the overdoped regime of high-$T_{\rm c}$ cuprates, it would then be possible to further optimize the $T_{\rm c}$ of RP bilayer nickelates provided a suitable tuning strategy.
Methods
Thin film growth and characterisation
LPNO thin films of approximately 5-nm thickness with a single-unit-cell capping layer of SrTiO3 were grown on SrLaAlO4(001) substrates using pulsed-layer deposition, as detailed in previous work [15]. Gold electrodes of 40-nm thickness with Hall-bar (or van der Pauw) geometry were deposited prior to ozone annealing in a tube furnace. Following transport characterisation measurements conducted using 4He cryostats, the samples were kept cold during storage at liquid nitrogen temperature and transported to high-field facilities at dry-ice temperature.
Magnetotransport measurements
Ultrasonically bonded aluminum wires were connected to the measurement probe using 25 $\mu$m-diameter gold wires and room-temperature silver paint (DuPont 4929N). Samples were exposed to ambient conditions for less than two hours between taken out of cryogenic storage and loaded into measurement cryostat. Electrical transport measurements with an AC excitation current of 10 $\mu$A were performed in static magnetic fields up to 13 T using a Physical Properties Measurement System by Quantum Design Inc. and in pulsed magnetic fields up to 53.4 T and 64.2 T at the International MegaGauss Science Laboratory in Institute for Solid State Physics, University of Tokyo and Dresden High Magnetic Field Laboratory (HLD-EMFL), Helmholtz-Zentrum Dresden-Rossendorf, respectively.
Extrapolation of zero-field resistivity below $T_{\rm c}$
The magnetoresistivity data at magnetic fields above which $\rho_{\rm xx}(H)$ shows a positive curvature is fitted to: $\rho(H,T)=\rho(0,T)+\beta H^{2}$ for extracting the zero-field resistivity $\rho_{\rm xx}(0,T)$ below $T_{\rm c}$. We further fine-tune the extracted $\rho_{\rm xx}(0,T)$ values to achieve a scaling collapse of the Kohler plot for $\rho_{\rm xx}(H,T)$ measured over the entire temperature range. The adjustment of $\rho_{\rm xx}(0,T)$ required for Kohler scaling shown in Fig. 2b is less than 0.5 $\mu\Omega\leavevmode\nobreak$cm for all temperatures.
Extraction of normal-state parameters
We use the 2D free electron model to extract the normal-state electronic structure parameters. The 2D carrier density ($n_{\rm 2D}$) is calculated using:
$n_{\rm 2D}=n_{\rm H}\times d$ (5)
where $d$ is the distance between Ni-O bilayers in adjacent unit cells. The Fermi wavevector, Fermi velocity, Fermi energy, and effective Fermi temperature are then calculated using:
$k_{\rm F}=(2\pi n_{\rm 2D})^{1/2},$ (6)
$v_{\rm F}=\hbar k_{\rm F}/m^{*},$ (7)
$E_{\rm F}=\hbar^{2}k_{\rm F}^{2}/(2m^{*}),$ (8)
$T_{\rm F}^{*}=\frac{\hbar^{2}\pi n_{\rm 2D}}{m^{*}k_{\rm B}},$ (9)
where $\hbar$ and $k_{\rm B}$ are the reduced Planck and Boltzmann constant, respectively, and $m^{*}/m_{e}$ as estimated using the measured $A_{2}$ coefficient and the empirical Kadowaki-Woods ratio of $10^{5}\leavevmode\nobreak\ \mu\Omega\leavevmode\nobreak\ \textrm{cm}\leavevmode\nobreak\ \textrm{mJ}^{-2}\leavevmode\nobreak\ \textrm{mol}^{-2}\leavevmode\nobreak\ \textrm{K}^{-6}$.
Data availability
The source data used to create the main figures are presented with this paper. Additional data are available from the corresponding authors upon request.
Author contributions
Y.-T.H. and Y.Y. conceived the project. Y.Y., Y.-T.H. and H.Y.H. designed the project. Y.L., Y.T. and Y.Y. synthesised and characterised the bilayer nickelate thin films under H.Y.H.’s supervision. Y.-T.H., Y.K., T.K. and V.S. performed the high-field magnetotransport measurements. Y.-T.H. analysed the data and wrote the manuscript with input from all authors.
Acknowledgements
We thank Toni Helm for experimental support with pulsed-field measurements and Bai Yang Wang for stimulating discussions. We also thank Mark Gibson and Jiarui Li for assistance in sample preparations. This work is supported by the Yushan Fellow Program (MOE-112-YSFMS-0002-002-P1) and the Center for Quantum Science and Technology (CQST) within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE), Taiwan, and by the National Science and Technology Council, Taiwan (NSTC 113-2112-M-008-044-MY3) (Y.-T.H). This work is also supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division under contract no. DE-AC02-76SF00515 (synthesis and transport measurements) as well as SuperC and the Kavli Foundation (in situ transport) (Y.L., Y.T., Y.Y., H.Y.H.), and by Grants-in-Aid for Scientific Research (KAKENHI, grant numbers 22H00104 and 25H00600) from the Japanese Society for the Promotion of Science (JSPS) (Y.K., V.S.). We further acknowledge support of the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL) as well as support under the European Union’s Horizon 2020 research and innovation programme through the ISABEL project (No. 871106).
Competing interests
The authors declare no competing interests.